A spinner with 9 equal sections is numbered 1 through 9. The probability of spinning an even number is 4/9 . What is the probability of not spinning an even number?
5/9 there are 4 even numbers between 1 and 9 and there are 5 odd numbers between 1 and 9 so if the probability of spinning an even number is 4/9 then the probability of spinning an odd number is 5/9 because 9-5 is 4 and 9-4 is 5
The probability of not spinning an even number is \(\mathbf{\dfrac{5}{9}}\)
Probability in mathematics is the possibility or likelihood of an event to occur. It is a measure of the number of required outcomes divided by the number of the total possible outcome of an event.
From the given information:
P(B) = 1 - P(A)
\(\mathbf{P(B) = 1 - \dfrac{4}{9}}\)
\(\mathbf{P(B) = \dfrac{5}{9}}\)
Learn more about probability here:
https://brainly.com/question/24756209